Question 10of 12 Step 4 of 10 0023 41 A manager for an insurance company beleves that customers have the folowr preferences for life insurance products: 20 % preter whole Life. 30 % prefer unversal Life, and 50 % prefer ufe Annuities. The results of a survey of 271 customers were tabulated. Is it possible to refute the saies managers claimed proportions of customers who prefer each product using the data? Product Number whole 134 Unversal111 Annuities 26 Copy Data Step 4 of 10 Find the expected value for the number of customers who prefer Whole Life. Round your answer to two decmai places Answenhaw to Ener) 2 Point E keypad Tables
Answer #1

Here, n = sample size = 271.

Thus, the expected value for the number of customers who prefer Whole Life = n * p1 = 271 * 0.2 = 54.20.